### v1

# MigGPT: Harnessing Large Language Models for Automated Migration of Out-of-Tree Linux Kernel Patches Across Versions

Abstract
Out-of-tree kernel patches are essential for adapting the Linux kernel to new hardware or enabling specific functionalities. Maintaining and updating these patches across different kernel versions demands significant effort from experienced engineers. Large language models (LLMs) have shown remarkable progress across various domains, suggesting their potential for automating out-of-tree kernel patch migration. However, our findings reveal that LLMs, while promising, struggle with incomplete code context understanding and inaccurate migration point identification. In this work, we propose MigGPT, a framework that employs a novel code fingerprint structure to retain code snippet information and incorporates three meticulously designed modules to improve the migration accuracy and efficiency of out-of-tree kernel patches. Furthermore, we establish a robust benchmark using real-world out-of-tree kernel patch projects to evaluate LLM capabilities. Evaluations show that MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 72.59% ($\uparrow 50.74\%$) for migration tasks.
Machine Learning, ICML
1 Introduction
The Linux kernel, a widely-used open-source operating system, is extensively applied across various domains (Tan et al., 2020; Lin et al., 2022; de Oliveira et al., 2023). Its adaptability and extensibility enable developers to create out-of-tree kernel patches that enhance performance (Kim et al., 2020; Adam, 2021) or security (Xu et al., 2023; Zhou et al., 2020), contributing to its widespread adoption. Out-of-tree kernel patches, such as RT-PREEMPT, AUFS, HAOC, Raspberry Pi kernel, and Open vSwitch, are modifications to the Linux kernel that are developed and maintained independently of the mainline source tree. Unlike in-tree patches, which are included in official kernel releases, out-of-tree patches address specific use cases or features not yet supported by the mainline kernel. As the Linux kernel evolves, these out-of-tree patches require ongoing maintenance to ensure compatibility with newer Linux kernel versions. As shown in Figure 1, the maintenance process involves utilizing the old out-of-tree kernel patch and analyzing the differences between the old and new Linux kernel versions to upgrade the patched kernel repository to the new version. This maintenance process demands specialized engineers and often takes weeks of intensive effort (Zhang et al., 2021).
Figure 1: MigGPT can assist in automating the version migration and maintenance of out-of-tree kernel patches of the Linux kernel. This saves on expert labor costs and reduces the development cycle.
Existing code migration technologies (Xing & Stroulia, 2007; Lamothe et al., 2022; Fazzini et al., 2019; Haryono et al., 2020; Thung et al., 2019; Ketkar et al., 2019; Rolim et al., 2017; Dilhara et al., 2023; Shi et al., 2022; Shariffdeen et al., 2021a, b; Yang et al., 2023) utilize static program analysis (Landi, 1992) to facilitate API cross-version maintenance or the backporting of CVE security patches. However, these methods only address a subset of scenarios in out-of-tree kernel patch migration. They rely on predefined migration rules, which are insufficient for handling comprehensive scenarios involving complex changes such as namespace modifications, invocation conflict resolution, and the integration of control and data flow dependencies.
With the substantial progress made by Large Language Models (LLMs) in understanding (Patel et al., 2023; Izadi et al., 2022; Kim et al., 2021) and generating code (Feng et al., 2020; Ahmed & Devanbu, 2022; Ouyang et al., 2023; Yan et al., 2023; Yang et al., 2024), there is a promising opportunity to leverage LLMs for the automated migration and maintenance of out-of-tree kernel patches. However, due to the inherent lack of determinism in LLMs when generating content, several challenges arise when directly employing these models to handle the migration and maintenance of out-of-tree kernel patches. These challenges include 1) interference from similar namespaces, 2) misalignment of code lines, 3) absence of associated code information, and 4) inaccuracies in locating migration points.
To address the challenges , we propose MigGPT, the first framework designed to assist humans in automating the migration and maintenance of out-of-tree kernel patches. MigGPT utilizes Code Fingerprint (CFP), a meticulously designed novel data structure to encapsulate the structural and critical information of code snippets throughout the migration process of out-of-tree kernel patches. With the assistance of CFP, MigGPT incorporates three core modules: the Retrieval Augmentation Module (addressing challenges 1 and challenges 3), the Retrieval Alignment Module (addressing challenge 2), and the Migration Enhancement Module (addressing challenge 4). Specifically, the Retrieval Augmentation Module supplies code snippet information via CFPs, mitigates interference from similar namespaces, and appends additional code snippets pertinent to migration. The Retrieval Alignment Module achieves alignment of the target code snippets through the first anchor line and the last anchor line within CFPs. The Migration Enhancement Module facilitates accurate and efficient migration by comparing CFPs to ascertain the number of migration points and their respective locations.
To evaluate the efficiency of MigGPT, we construct a robust benchmark that includes three real-world projects from the out-of-tree patch community of the Linux kernel. These projects comprise two different levels of migration examples, encompassing a variety of common migration types. With this benchmark, we evaluated MigGPT across diverse LLMs (GPT-3.5, GPT-4-turbo , DeepSeek-V2.5, DeepSeek-V3, and Llama-3.1) (OpenAI, 2023a, b; DeepSeek-AI et al., 2024a, b; Dubey et al., 2024) to validate its effectiveness and broad applicability. MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 72.59% ($\uparrow 50.74\%$) for migration tasks. Additionally, the average number of queries to LLMs is only 2.22($\uparrow 0.22$), indicating no substantial increase in computational overhead.
In summary, we make the following contributions:
  • We have developed a robust and migration benchmark, encompassing three real-world projects. To the best of our knowledge, this is the first benchmark for out-of-tree kernel patch migration that can assess performance across diverse migration tools, providing a valuable foundation for future research.
  • We propose CFP, a carefully designed data structure that encapsulates the structural and critical information of code snippets, providing essential migration context for LLMs. Based on this, we introduce MigGPT, a framework to assist humans in automating out-of-tree kernel patch migration and maintenance.
  • We conduct comprehensive experiments on both closed-source models (i.e. GPT-3.5 and GPT-4) and open-source models (i.e. DeepSeek-V2.5, DeepSeek-V3, and Llama-3.1). The results demonstrate that MigGPT achieved an average migration accuracy of 72.59%($\uparrow 52.74\%$), representing a significant improvement over vanilla LLMs.
2 Related Work
2.1 Code Migration Kernel Patch
Existing code migration (Dilhara et al., 2024; Pan et al., 2024) technologies primarily focus on API cross-version maintenance (Xing & Stroulia, 2007; Lamothe et al., 2022; Fazzini et al., 2019; Haryono et al., 2020; Thung et al., 2019; Ketkar et al., 2019; Rolim et al., 2017; Dilhara et al., 2023) and the backporting of CVE security patches (Shi et al., 2022; Shariffdeen et al., 2021a, b; Yang et al., 2023). Research on API cross-version maintenance employs static analysis (Landi, 1992) to detect deprecated API patterns and migrate code using transformation rules. However, these methods only partially address out-of-tree kernel patch migration due to the tight coupling between kernel and patch code. Similarly, CVE security patch backporting techniques identify vulnerability patterns and apply predefined rules but fail to manage complex changes such as namespace modifications, invocation conflicts, and control/data flow dependencies, limiting their effectiveness in comprehensive migration scenarios.
2.2 LLMs for Coding
In recent years, LLMs (Chen et al., 2021; Fried et al., 2023; Rozière et al., 2023; Le et al., 2022; Nijkamp et al., 2023; Li et al., 2023; OpenAI, 2023b) have achieved remarkable progress in various natural language processing tasks. Initially focused on natural language understanding and generation, the adaptability of LLMs has expanded to the field of software engineering, where they can be fine-tuned to perform programming tasks such as code completion (Patel et al., 2023; Izadi et al., 2022; Kim et al., 2021), code search (Feng et al., 2020), code summarization (Ahmed & Devanbu, 2022), code generation (Ouyang et al., 2023), and even complex code repair (Fu et al., 2022; Jesse et al., 2023). This inspires us to apply LLMs to the migration and maintenance of out-of-tree kernel patches. To the best of our knowledge, MigGPT is the first work to apply LLMs to this task, paving the way for subsequent research.
3 Preliminaries
Out-of-Tree Kernel Patches
Out-of-tree kernel patches are modifications outside the mainline Linux kernel, often created by third-party developers or organizations to add features, fix issues, or optimize performance for specific hardware. They provide quick access to new functionalities or hardware support, catering to specialized needs. Examples include RT-PREEMPT and AUFS for feature enhancements, HAOC for security improvements, and Raspberry Pi kernel and Open vSwitch for platform-specific optimizations. An example of out-of-tree kernel patch migration is provided in App. B.
Motivation.
Such out-of-tree kernel patches lack official support and require manual maintenance to ensure compatibility with future kernel versions, often demanding weeks of effort from specialized engineers. Automating this process could significantly reduce reliance on manual labor, thereby conserving substantial human resources and accelerating development cycles.
Problem Formulation.
Out-of-tree kernel patches lack official support and require manual maintenance to ensure compatibility with future Linux kernel versions. Let $R$ denote a Linux kernel repository, where $s\in R$ represents a code snippet within the repository. The older version of Linux kernel is $R_{\text{old}}$, and after applying an out-of-tree patch, it becomes $R^{\prime}_{\text{old}}$. When the kernel advances to a new version $R_{\text{new}}$, the migration problem is to construct a function $M:R^{\prime}_{\text{old}}\to R^{\prime}_{\text{new}}$ where $\forall s\in R^{\prime}_{\text{old}},\exists M(s)\in R^{\prime}_{\text{new}}$ s.t.$\forall x\in\text{Inputs},\text{Execute}(R^{\prime}_{\text{old}},x)=\text{Execute}(R^{\prime}_{\text{new}},x)$.
4 Migration Benchmark
4.1 Migration Types
We can obtain the code snippets $s_{\text{old}}\in R_{\text{old}}$ and $s^{\prime}_{\text{old}}\in R^{\prime}_{\text{old}}$ at the same location in the repository before and after applying the out-of-tree kernel patch, with the differences represented by $\Delta$. As $R_{\text{old}}$ is updated to a new version of the Linux kernel $R_{\text{new}}$, we need to locate the corresponding code snippet $s_{\text{new}}\in R_{\text{new}}$ in the new version of the Linux kernel to obtain the difference information during the kernel update. The differences between $s_{\text{old}}$ and $s_{\text{new}}$ are denoted as $\Sigma$. Subsequently, by utilizing the information from $\Delta$ and $\Sigma$, we complete the migration task to obtain the new version of the out-of-tree kernel patch code snippet $s^{\prime}_{\text{new}}$. Finally, these code snippets are integrated to form the new version of the out-of-tree kernel patch code repository $R^{\prime}_{\text{new}}$.
Considering the states of $\Delta$ and $\Sigma$, we can categorize the migration types into two classes:
Type 1: This type of migration example satisfies $\Delta\neq\varnothing,\Sigma\neq\varnothing$, $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle=0$. This indicates that both the out-of-tree kernel patch and the new version of the Linux kernel have modified the code snippet, and their changes do not affect the same lines of code, meaning the modifications do not overlap or conflict with each other.
Type 2: In contrast, this type satisfies $\Delta\neq\varnothing,\Sigma\neq\varnothing$, $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$, indicating that their modifications overlap on the same lines of code, leading to conflicts.
The remaining cases, $\Delta=\varnothing$ and $\Delta\neq\varnothing,\Sigma=\varnothing$, signify no code modification in the out-of-tree kernel patch and no changes in the new kernel version, respectively. Due to their simplicity and straightforward migration, they are excluded from our benchmark.
Table 1: Formalization, Counts of the Two Types of Migration Example. Other examples are too simple to necessitate resolution. More detailed examples are in the App. A.2.
Class Formalization Number Type 1 $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ 80 $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle=0$ (59.3%) Type 2 $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ 55 $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$ (40.7%) Others $\Delta=\varnothing$ or Too simple $\Delta\neq\varnothing,\Sigma=\varnothing$ to resolve
4.2 Benchmark Design
We have built a robust migration testing benchmark using out-of-tree kernel patches from real-world projects, specifically focusing on three open-source initiatives: RT-PREEMPT (Kernel, 2013), Raspberry Pi Linux (Pi, 2018), and HAOC (HAOC, 2024)  [^1] Even with knowledge of the code in these out-of-tree kernel patches, LLMs still struggle to accomplish migration and maintenance tasks.. RT-PREEMPT enhances the Linux kernel’s real-time performance for timing-sensitive applications like industrial control and robotics, while Raspberry Pi Linux offers a lightweight kernel optimized for embedded systems. HAOC improves kernel security through a ”dual-kernel” design, enhancing code behavior, data access, and permission management. We collect code from these projects across Linux kernel versions 4.19, 5.4, 5.10, and 6.6 for our benchmark.
Guided by the experience of manually completing the task, we divide the migration task into two steps: 1) Identifying the migration location, i.e., finding $s_{\text{new}}$. 2) Completing the migration to obtain $s^{\prime}_{\text{new}}$. In this case, firstly, we use the diff command to obtain the code snippets $s_{\text{old}}$ and $s^{\prime}_{\text{old}}$ from files with the same name in the code repository. Subsequently, by matching filenames, we locate the file in code repository $R_{\text{new}}$ that contains the target new version code snippet $s_{\text{new}}$. Finally, we gather the ground truth (results manually completed by humans) $\hat{s}_{\text{new}}$ and $\hat{s}^{\prime}_{\text{new}}$. Specifically, our benchmark includes a quintuple $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}},\hat{s}_{\text{new}},\hat{s}^{\prime}_{\text{new}})$ for each migration example. After filtering out invalid differences (such as spaces, blank lines, file deletions, etc.), we randomly collected 135 migration examples, comprising 80 Type 1 and 55 Type 2, as detailed in Table 1.
Figure 2: Overview of MigGPT. MigGPT employs a code fingerprint (CFP) structure to retain code snippet information, enhanced by three modules to improve migration accuracy and efficiency. The migration process involves two steps: 1) locating the migration position in $\text{file}_{\text{new}}$ to find $s_{\text{new}}$, and 2) completing the migration to obtain $s^{\prime}_{\text{new}}$.
5 MigGPT
We first outline the challenges faced when utilizing vanilla LLMs for the migration of out-of-tree kernel patches (Section 5.1), and then discuss how MigGPT effectively addresses these challenges (Sections 5.2 to 5.7).
5.1 Challenges
Through analyzing LLM behavior and results, we identify key challenges hindering their success in out-of-tree kernel patch migration:
Challenge 1 (Namespace Interference): When identifying the code snippet $s_{\text{new}}$ in $\text{file}_{\text{new}}$ for migration, retrieval errors can occur. LLMs often struggle to locate function definitions within $s_{\text{new}}$ due to interference from similarly named functions, leading to inaccuracies that affect subsequent migration stages. An example is provided in App. C.1.
Challenge 2 (Misalignment): This challenge occurs when retrieving $s_{\text{new}}$ from $\text{file}_{\text{new}}$. Due to the inherent randomness in LLM-generated responses, discrepancies often arise between the start and end lines of $s_{\text{new}}$ identified by the LLM and those retrieved by human developers ($\hat{s}_{\text{new}}$). This misalignment can result in missing or extraneous lines, significantly compromising migration outcomes where precise code segment boundaries are critical. An example is provided in App. C.2.
Challenge 3 (Missing Associated Fragments): This challenge occurs when retrieving $s_{\text{new}}$ from $\text{file}_{\text{new}}$. During Linux kernel upgrades, code blocks from older versions may be split into fragments in the new version for standardization or reuse. LLMs often fail to identify and retrieve all these fragments, leading to incomplete $s_{\text{new}}$. This results in errors during out-of-tree kernel patch migration due to missing code segments. An example is provided in App. C.3.
Challenge 4 (Ambiguous Migration Points): This challenge arises during the migration of $s_{\text{new}}$ to $s^{\prime}_{\text{new}}$. Although the information provided by $s_{\text{old}}$ and $s^{\prime}_{\text{old}}$ is sufficient to accurately infer the migration point, LLMs frequently fail to precisely identify these points. This ambiguity results in errors when determining the correct location for migration. An example is provided in App. C.4.
Overall, LLMs require migration-relevant code structure information and code scope constraints to more effectively migrate and maintain out-of-tree kernel patches.
5.2 Overview
To this end, we propose MigGPT, a framework combining traditional program analysis with LLMs to facilitate out-of-tree kernel patch migration across Linux versions. As outlined in Section 4, MigGPT works in two stages: identifying target code snippets in the new version and migrating the out-of-tree patch. Figure 2 shows its three core modules: the Retrieval Augmentation Module (addressing Challenges 1 and 3), the Retrieval Alignment Module (addressing Challenge 2), and the Migration Enhancement Module (addressing Challenge 4). Each module uses a code fingerprint structure, which encodes the structural features of code snippets, to enhance LLM performance and migration accuracy, tackling the challenges discussed earlier.
5.3 Code Fingerprint
To address the challenges LLMs face in migrating out-of-tree kernel patches across Linux kernel versions, a detailed analysis of code snippet structure is essential to identify migration-related code structure information and code scope constraints. While tools like Abstract Syntax Tree (AST) are useful for structural analysis, they have limitations: 1) Inability to process incomplete or non-compilable code snippets due to tight integration with the compilation process. 2) The mismatch between excessive structural details (AST tools provide a plethora of information irrelevant to patch migration) and the absence of critical information (such as comments and inline assembly), which is essential for maintaining and updating kernel patches  [^2] Inline assembly is widely used in the Linux kernel, and comments are crucial for future module development, as their omission would hinder subsequent modifications.. Focusing on key statements, such as migration points and alignment positions, while preserving essential elements like comments and inline assembly, can enhance efficiency and reduce overhead in the migration process.
To address the limitations of traditional code structure analysis, we propose Code Fingerprint (CFP), a lightweight sequential data structure for analyzing code snippets. CFP records both the content and positional information for each line of statements, encompassing all C language statements, including comments and inline assembly (a detailed example is provided in App. D.1). As shown in Figure 4, CFP focuses on recording function definitions and function calls, which are crucial for addressing challenges 1 and 3, as detailed in Section 5.4. Additionally, its linear structure facilitates accurate positioning for insertion, deletion, and other update operations, tackling challenges 2 and 4, further explained in Sections 5.5 and  5.6. Overall, CFP offers three key advantages: 1) effective processing of incomplete code snippets, 2) preservation of critical information such as comments and inline assembly, which are vital for out-of-tree kernel patch migration, and 3) a streamlined design that focuses on essential statements, improving migration efficiency and reducing overhead. By minimizing unnecessary processing while ensuring relevance, CFP provides a targeted solution for migrating out-of-tree kernel patches across Linux kernel versions.
Figure 3: A C code snippet containing inline assembly statements and comment annotations.
Figure 4: Compared to AST, CFP extracts key code structures, and its linear representation enables clearer localization of code modification points.
5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function signatures when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the namespace (“Namespace Prompt”), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2 Figure 8.
On the other hand, challenge 3 highlights that during the migration update of out-of-tree kernel patches by LLMs, there is an issue with missing associated functions. For the LLM’s temporary retrieval result $s_{\text{tmp}}$, we utilize the code fingerprint structures $\text{CFP}_{\text{tmp}}$ and $\text{CFP}_{\text{old}}$ of $s_{\text{tmp}}$ and $s_{\text{old}}$, respectively, and extract from them the sets of internally called associated functions, denoted as $\mathcal{F}_{\text{tmp}}$ and $\mathcal{F}_{\text{old}}$. Then, using string matching techniques, we retrieve from $\text{file}_{\text{new}}$ the code snippets corresponding to the associated function calls $Funccall|$ that satisfy $Funccall|\in\mathcal{F}_{\text{tmp}}\setminus\mathcal{F}_{\text{old}}$. Ultimately, these associated function code snippets are combined with $s_{\text{tmp}}$ to form the complete code snippet $s_{\text{new}}$. An example is provided in the App. D.2 Figure 10.
5.5 Retrieval Alignment Module
The retrieval alignment module is devised to tackle challenge 2, which was encountered during the migration update of out-of-tree kernel patches by LLMs. Challenge 2 indicates that when the LLM retrieves the target code snippet $s_{\text{new}}$ from the new version of the Linux kernel file $\text{file}_{\text{new}}$, there can be a mismatch between the start and end lines of $s_{\text{new}}$. To address this issue, we need to leverage the information from the first and last lines of the old version code snippet $s_{\text{old}}$ to aid in the localization during the retrieval of $s_{\text{new}}$. As illustrated in Figure 2, we utilize the code fingerprint structure $\text{CFP}_{\text{old}}$ of $s_{\text{old}}$. By taking advantage of its linear structure, we obtain the CFP statements for the first and last lines. These statements are used to construct an “Alignment Prompt”, which describes the information of the first and last lines and is included as part of the input to the LLM. This prompt guides the LLM in performing the retrieval task better by accurately identifying the boundaries of the target code snippet.
Algorithm 1 Retrieval of the target code snippet $s_{\text{new}}$
1:  Input: $\left(s_{\text{old}},\text{file}_{\text{new}}\right)$, $\operatorname{LLM}$, and maximum query count $m$
2:  Output: $s_{\text{new}}$
3:  Generating $\text{CFP}_{\text{old}}$ form $s_{\text{old}}$
4:  Preparing $RetrievalTaskPrompt$
5:  Preparing $ExpertPersonaPrompt$
6:  $\mathcal{S}\leftarrow\operatorname{Extractsignature}(\text{CFP}_{\text{old}})$
7:  $NamespacePrompt\leftarrow\operatorname{Prompt}(\mathcal{S})$
8:  $\mathcal{A}\leftarrow\operatorname{Extractanchor}(\text{CFP}_{\text{old}})$
9:  $AlignmentPrompt\leftarrow\operatorname{Prompt}(\mathcal{A})$
10:  $RetrievalPrompt\leftarrow$
11:   $+RetrievalTaskPrompt+NamespacePrompt$
12:   $+AlignmentPrompt+ExpertPersonaPrompt$
13:  while $q<=m$ do
14:     $s_{tmp}\leftarrow\operatorname{LLM}(RetrievalPrompt,s_{\text{old}},\text{file}_{\text{new}})$
15:     Generating $\text{CFP}_{\text{tmp}}$ from $s_{tmp}$
16:     if $\operatorname{find}(\mathcal{S},s_{tmp})$ then
17:        break
18:     end if
19:     $q\leftarrow q+1$
20:  end while
21:  $\mathcal{F}_{\text{old}}\leftarrow\operatorname{Funccall}(\text{CFP}_{\text{old}})$
22:  $\mathcal{F}_{\text{tmp}}\leftarrow\operatorname{Funccall}(\text{CFP}_{\text{tmp}})$
23:  $s_{\text{new}}\leftarrow s_{\text{tmp}}+\operatorname{FindCode}(\mathcal{F}_{\text{tmp}}\setminus\mathcal{F}_{\text{new}},\text{file}_{\text{new}})$
24:  return $s_{\text{new}}$
Algorithm 2 Migration of code snippet $s^{\prime}_{\text{new}}$
1:  Input: $\left(s_{\text{old}},s^{\prime}_{\text{old}},s_{\text{new}}\right)$ and $\operatorname{LLM}$
2:  Output: $s^{\prime}_{\text{new}}$
3:  Generating $\text{CFP}_{\text{old}}$, $\text{CFP}^{\prime}_{\text{old}}$ form $s_{\text{old}}$ and $s^{\prime}_{\text{old}}$
4:  Preparing $MigrationTaskPrompt$
5:  Preparing $ExpertPersonaPrompt$
6:  $\mathcal{P}\leftarrow\operatorname{PinpointMigrationLocation}(\text{CFP}_{\text{old}},\text{CFP}^{\prime}_{\text{old}})$
7:  $LocationPrompt\leftarrow\operatorname{Prompt}(\mathcal{P})$
8:  $MigrationPrompt\leftarrow+MigrationTaskPrompt$
9:   $+LocationPrompt+ExpertPersonaPrompt$
10:  $s^{\prime}_{\text{new}}\leftarrow\operatorname{LLM}(LocationPrompt,s_{\text{old}},s^{\prime}_{\text{old}},s_{\text{new}})$
11:  return $s^{\prime}_{\text{new}}$
5.6 Migration Augmentation Module
The migration augmentation module is primarily designed to address challenge 4 encountered by LLMs during the migration of new-version Linux kernel code snippets $s_{\text{new}}$ into the final updated out-of-tree kernel patch $s^{\prime}_{\text{new}}$. In challenge 4, LLMs often struggle to accurately identify the number and location of migration points, leading to errors in the final migrated $s^{\prime}_{\text{new}}$. As illustrated in Figure 2, to tackle this challenge, we leverage information from the old version code snippet $s_{\text{old}}$ and its modified counterpart $s^{\prime}_{\text{old}}$ to determine the number and location of modifications made to the out-of-tree kernel patch. This information is used to construct a “Location Prompt” that assists the LLM in accurately identifying the number and location of migration points. An example is provided in the App. D.3.
5.7 Implementation
With the critical code information provided by CFP, we can leverage the Retrieval Augmentation Module and the Retrieval Alignment Module to assist LLMs in more effectively identifying target kernel code snippets $s_{\text{new}}$. Subsequently, with the aid of the Migration Augmentation Module, we facilitate the migration to generate the final code snippet $s^{\prime}_{\text{new}}$. All the prompts of MigGPT are provided in App. D.4.
As illustrated in Algorithm 1, we first need to retrieve the target code snippet $s_{\text{new}}$ from $\text{file}_{new}$. Specifically, using the information contained within $\text{CFP}_{old}$, we can extract a set of critical function signatures $\mathcal{S}$ and a set of key anchor statements $\mathcal{A}$. With this information, we construct the $NamespacePrompt$ and $AlignmentPrompt$, ultimately forming the complete $RetrievalPrompt$. We then query LLMs using the $RetrievalPrompt$ to obtain an initial result $s_{\text{tmp}}$. We check if $s_{\text{tmp}}$ contains items from the target function signature set $\mathcal{S}$. If not, we repeatedly query the LLMs using the $RetrievalPrompt$ (up to $m$ times). If it does contain items from $\mathcal{S}$, we use $\text{CFP}_{old}$ and $\text{CFP}_{tmp}$ to extract newly appeared called functions within $s_{\text{tmp}}$ and retrieve the code snippets where these called functions are defined from $\text{file}_{new}$ as additional supplementary information for $s_{\text{tmp}}$. Finally, we concatenate these two parts of the code snippets to obtain $s_{\text{new}}$.
After obtaining $s_{\text{new}}$, we proceed to migrate it to achieve $s^{\prime}_{\text{new}}$. As shown in Algorithm 2, we utilize the differences between $\text{CFP}_{old}$ and $\text{CFP}^{\prime}_{old}$ to extract the number and positions of migration points and generate the $LocationPrompt$. Further, we formulate the $MigrationPrompt$ and query the LLM to obtain the migrated out-of-tree kernel patch code snippet $s^{\prime}_{\text{new}}$.
6 Evaluation
In this section, we assess the performance of MigGPT, focusing on the following three research questions:
RQ1 (Performance): How does the performance of MigGPT compare with that of vanilla LLM models?
RQ2 (Ablation): How does each module within MigGPT contribute to the overall migration performance?
RQ3 (Failure Analysis): How much modification is required for MigGPT’s failed example to align with human-level performance in out-of-tree patch migration tasks?
Table 2: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs in retrieving target code snippets.
LLM Method Type 1 (80) Type 2 (55) All (135) Average Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match Query Times GPT-3.5 Vanilla 20.00% 33.75% 26.25% 20.00% 25.45% 27.27% 20.00% 30.37% 26.67% 1.00 MigGPT 68.75% 68.75% 71.25% 61.82% 54.55% 70.91% 65.93% 62.96% 71.11% 1.28 GPT-4-turbo Vanilla 60.00% 67.50% 65.00% 69.09% 76.36% 78.18% 63.70% 71.11 70.37% 1.00 MigGPT 91.25% 95.00% 96.25% 81.82% 83.64% 89.09% 87.41% 90.37% 93.33% 1.16 DeepSeek-V2.5 Vanilla 61.25% 66.25% 62.50% 65.45% 69.09% 67.27% 62.96% 67.40% 64.44% 1.00 MigGPT 95.00% 97.5% 96.25% 87.27% 90.90% 90.90% 91.85% 94.81% 94.07% 1.22 DeepSeek-V3 Vanilla 68.75% 71.25% 72.50% 74.55% 78.18% 78.18% 71.11% 74.07% 74.81% 1.00 MigGPT 92.50% 93.75% 95.00% 85.45% 78.18% 89.09% 89.63% 87.41% 92.59% 1.22 Llama-3.1-70B Vanilla 58.75% 65.00% 63.75% 61.82% 72.73% 75.55% 60.00% 68.15% 68.15% 1.00 MigGPT 91.25% 92.50% 93.75% 80.00% 81.82% 81.82% 86.67% 88.15% 88.89% 1.24 Average Vanilla 53.75% 60.75% 58.00% 58.18% 64.36% 65.09% 55.56% 62.22% 60.89% 1.00 MigGPT 87.75% 89.50% 90.50% 79.27% 77.82% 84.36% 84.30% 84.74% 88.00% 1.23 $\uparrow$ +34.00% +28.75% +32.50% +21.09% +13.45% +19.27% +28.74% +22.52% +27.11% -
Table 3: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs in the migration task of target code snippets.
LLM Method Type 1 (80) Type 2 (55) All (135) Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match GPT-3.5 Vanilla 7.50% 5.00% 8.75% 3.64% 3.64% 5.45% 5.93% 4.44% 7.41% MigGPT 37.50% 46.26% 47.50% 38.18% 41.82% 61.82% 37.78% 44.44% 53.33% GPT-4-turbo Vanilla 15.00% 12.50% 18.75% 10.91% 30.91% 23.64% 13.33% 20.00% 20.74% MigGPT 68.75% 82.50% 85.00% 54.55% 76.36 69.09% 62.96% 80.00% 78.52% DeepSeek-V2.5 Vanilla 11.25% 16.25% 18.75% 9.09% 27.27% 21.82% 10.37% 20.74% 20.00% MigGPT 67.50% 80.00% 80.00% 56.36% 74.55% 70.91% 62.96% 77.78% 76.30% DeepSeek-V3 Vanilla 23.75% 37.50% 32.50% 34.55% 54.55% 49.09% 28.15% 44.44% 39.26% MigGPT 81.25% 88.75% 87.50% 65.45% 78.18% 74.55% 74.81% 84.44% 82.22% Llama-3.1-70B Vanilla 3.75% 16.25% 18.75% 7.27% 27.27% 23.64% 5.19% 20.74% 20.74% MigGPT 62.50% 80.00% 81.25% 47.27% 67.27% 72.73% 56.30% 74.81% 77.78% Average Vanilla 12.25% 17.50% 19.50% 13.09% 28.73% 24.73% 12.59% 22.07% 21.63% MigGPT 63.50% 75.50% 76.25% 52.36% 67.64% 69.82% 58.96% 72.30% 73.63% $\uparrow$ +51.25% +58.00% +56.75% +39.27% +38.91% +45.09% +46.37% +50.22% +52.00%
6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark (Shariffdeen et al., 2021a), which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 (OpenAI, 2023a), GPT-4-turbo (OpenAI, 2023b), DeepSeek-V2.5 (DeepSeek-AI et al., 2024a), DeepSeek-V3 (DeepSeek-AI et al., 2024b), and Llama-3.1-70B-Instruct (Dubey et al., 2024), as they are widely recognized for their advanced capabilities and performance in code-related tasks, along with patch backporting tools like Patch (GNU, 2020), SyDIT (Meng et al., 2011), and FixMorph (Shariffdeen et al., 2021a). Evaluation metrics include “best match” (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match” (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. E.2(Ren et al., 2020), and “human match” (developer-judged functional equivalence). The hyperparameter $m$ is set to 3.
For each migration sample $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}},\hat{s}_{\text{new}},\hat{s}^{\prime}_{\text{new}})$ in our benchmark, we evaluate vanilla LLMs using two distinct strategies: One-step Strategy: The LLM directly generates the migrated code snippet $s^{\prime}_{\text{new}}$ by taking the triplet $(s_{\text{old}},s^{\prime}_{\text{old}},\text{file}_{\text{new}})$ as input. Two-step Strategy: The process is divided into two phases. First, the LLM identifies the corresponding new version code snippet $s_{\text{new}}$ using the pair $(s_{\text{old}},\text{file}_{\text{new}})$. Then, the LLM generates $s^{\prime}_{\text{new}}$ by taking the triplet $(s_{\text{old}},s^{\prime}_{\text{old}},s_{\text{new}})$ as input.
Figure 5: The semantic match accuracy of target code snippets retrieval task and target code snippets migration task across various LLMs. “One-step” indicates the direct utilization of an LLM to complete the migration task in a single step.
6.2 Performance Evaluation (RQ1)
MigGPT demonstrates exceptional capability in retrieving target code snippets.
As shown in Table 2, in the first step of the two-step strategy–retrieving target code snippets from new versions of the Linux kernel–MigGPT exhibits a significant advantage in the subtask of retrieving target code. Specifically, when paired with a high-performance LLM like GPT-4-turbo, MigGPT achieves a language matching precision of 96.25% for Type 1 samples, significantly outperforming standalone GPT-4-turbo (65.00%). Overall, MigGPT attains an average semantic matching precision of 83.89% across all sample types, marking a 23.15% relative improvement.
MigGPT demonstrates outstanding performance in generating migrated code snippets.
As shown in Table 3, in the second step of the two-step strategy—-generating code snippets for new versions of out-of-tree kernel patches—-MigGPT outperforms vanilla LLMs, achieving a 49.26% higher average migration semantic matching precision, a 220% relative improvement. Performance gains increase with the underlying LLM’s capabilities, reaching 80.00% with MigGPT-augmented GPT-4-turbo and 84.44% with MigGPT-augmented DeepSeek-V3.
The two-step strategy outperforms the one-step strategy.
We compared GPT-3.5 and GPT-4-turbo using both one-step and two-step strategies to investigate the impact of task complexity on migration performance. As illustrated in Figure 5, when vanilla LLMs are employed, the two-step strategy achieves an average migration accuracy of 12.22% across all sample types, representing an improvement of 8.89% over the one-step strategy’s accuracy of 3.33%.
MigGPT excels in CVE patch backporting task.
We also evaluate the performance of MigGPT in the context of CVE patch backporting. As illustrated in Table 4, MigGPT demonstrates superior performance compared to existing patch backporting methods.
Table 4: The semantic match accuracy of MigGPT compared to patch backporting methods.
Method Patch SyDIT FixMorph GPT-4-turbo DeepSeek-V3 vanilla MigGPT vanilla MigGPT Accuracy 36% 28% 75% 85% 91% 87% 92%
6.3 Ablation Study (RQ2)
We conduct an ablation study to evaluate the impact of the four units in MigGPT: CFP, Retrieval Augmentation Module, Retrieval Alignment Module, and Migration Augmentation Module (details in App. E.3). Figure 6 presents the outcomes of four tested variants on our benchmarks. Among these, MigGPT consistently outperforms the ablation baselines in both retrieval and migration tasks. Meanwhile, we perform an ablation study on hyper-parameter $m$ in Algorithm 1 with MigGPT-GPT-4-turbo, which controls the total query time of the Retrieval Augmentation Module. As shown in Figure 6, $m=3$ is suitable for both Type 1 and Type 2 examples.
Figure 6: Left: The accuracy of different variants of MigGPT. Right: The best match retrieval accuracy of different $m$.
6.4 Failure Analysis (RQ3)
We evaluate MigGPT’s robustness by analyzing failed migration cases (not human-matched) across various samples, measuring line edit distances (insertions, deletions, modifications) between MigGPT’s incorrect outputs and human-corrected results (see App. E.1). As shown in Table 5, 41% of MigGPT’s errors require fewer than three lines of modification to align with correct results, demonstrating its potential to aid in out-of-tree kernel patch migration.
Table 5: The number of line edit distances between failed cases and human-migrated outcomes in migration task. “$3\leq\text{dis}<6$” denotes a line edit distance of at least 3 but less than 6.
LLM Type $\text{dis}<3$ $3\leq\text{dis}<6$ $6\leq\text{dis}<9$ $9\leq\text{dis}$ All GPT-3.5 Type 1 13 9 8 12 42 Type 2 8 4 2 7 21 GPT-4-turbo Type 1 5 1 3 3 12 Type 2 9 1 0 7 17 DeepSeek-V2.5 Type 1 10 2 3 1 16 Type 2 8 1 3 4 16 DeepSeek-V3 Type 1 3 2 3 2 10 Type 2 5 4 1 4 14
7 Conclusion
This study explores the migration of out-of-tree kernel patches in the Linux kernel across various versions. Our proposed benchmark reveals that LLMs struggle with incomplete code context understanding and inaccurate migration point identification. To address these issues, we propose MigGPT, an automated tool for migrating Linux kernel downstream modules. Our evaluation highlights MigGPT’s effectiveness and potential to advance this field.
Impact Statement
This work advances the field of automated software maintenance by introducing MigGPT, a framework that leverages LLMs to automate the migration and maintenance of out-of-tree Linux kernel patches. By reducing the manual effort and costs associated with these tasks, our research has the potential to improve the efficiency and reliability of software systems, benefiting industries that rely on stable and up-to-date infrastructure.
However, the adoption of such automation tools also raises ethical considerations. For example, automating tasks traditionally performed by specialized engineers may impact job roles, necessitating workforce adaptation. Additionally, the reliance on LLMs for critical maintenance tasks requires rigorous validation to ensure accuracy and avoid potential risks to system stability and security.
While our primary focus is on technical advancements, we acknowledge the broader societal implications of automating complex engineering processes. This study lays the foundation for future research and encourages ongoing discussions on the responsible use of AI in software maintenance, balancing innovation with ethical considerations.
Appendix A Benchmark
A.1 Collection
The migration examples in our benchmark are derived from three open-source out-of-tree kernel patch projects: RT-PREEMPT (Kernel, 2013), HAOC (HAOC, 2024) and Raspberry Pi kernel (Pi, 2018). Notably, RT-PREEMPT’s latest version has been integrated into the mainline Linux kernel for maintenance and no longer exists as an out-of-tree kernel patch. However, this does not impede our utilization of it for research on automated migration and maintenance of out-of-tree kernel patches.
A.2 Examples of Benchmark
As shown in Table 1, we categorized these samples based on the difficulty of migration into two classes:
Type 1: This type of migration example satisfies $\Delta\neq\varnothing,\Sigma\neq\varnothing$, $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle=0$. This indicates that both the out-of-tree kernel patch and the new version of the Linux kernel have modified the code snippet, and their changes do not affect the same lines of code, meaning the modifications do not overlap or conflict with each other. As shown in Table 1 for example, $s^{\prime}_{\text{old}}$ introduces additional lines of code to the function definition of hisilicon_1980005_enable in $s_{\text{old}}$. Conversely, $s_{\text{new}}$ both adds and removes certain lines of code within the same function definition in $s_{\text{old}}$. However, it is important to note that these modifications do not occur on the same lines of code.
Type 2: This type of migration example satisfies $\Delta\neq\varnothing,\Sigma\neq\varnothing$, $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$. This indicates that both the out-of-tree kernel patch and the new version of the Linux kernel have modified the code snippet, and their changes affect the same lines of code, resulting in overlapping modifications that conflict with each other. As illustrated in Table 1, for instance, $s^{\prime}_{\text{old}}$ introduces additional lines of code to the function definition of ptep_get_and_clear in $s_{\text{old}}$. However, $s_{\text{new}}$ refactors the same function definition into two separate function definitions, resulting in overlapping modifications that conflict with each other.
Table 6: Formalization, Counts, and Examples of the Three Types of Migration Example.
Class Type 1 Type 2 Formalization $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$ Number 80 (48.2%) 86 (51.8%) $s_{\text{old}}$ vs $s^{\prime}_{\text{old}}$ [Uncaptioned image] [Uncaptioned image] $s_{\text{old}}$ vs $s_{\text{new}}$ [Uncaptioned image] [Uncaptioned image]
Appendix B Examples of Out-of-tree Kernel Patch Migration
As shown in Figure 7, the migration maintenance of an out-of-tree kernel patch requires integrating the modifications from the old version out-of-tree kernel patch and the modifications from the new version Linux kernel to ultimately complete the code snippet for the new version out-of-tree kernel patch.
Figure 7: (a) Old version Linux kernel code snippet, with the green section indicating modifications from the old version out-of-tree kernel patch; (b) Old version Linux kernel code snippet, with the red and green sections indicating modifications for the new Linux version kernel; (c) New Linux version kernel code snippet, with the green section indicating modifications from the new version out-of-tree kernel patch.
Appendix C Examples of Each Challenge
C.1 Challenge 1
Figure 8: A migration case for challenge 1. The green code denotes modifications originating from the out-of-tree kernel patches.
In the migration case shown in Figure 8, we need to locate the target code snippet $s_{\text{new}}$, which defines the function __pte_free_tlb, within the code file $\text{file}_{\text{new}}$ of the new Linux kernel version. However, the new version file also contains a code snippet __pmd_free_tlb that closely resembles the target code snippet __pte_free_tlb. When LLMs attempt to locate the function __pte_free_tlb in $\text{file}_{\text{new}}$, they erroneously retrieve the similar function __pmd_free_tlb. This misidentification leads to errors during the migration of the out-of-tree kernel patch code. This issue highlights the challenges faced by LLMs in distinguishing between similar elements within codebases, indicating a need for improved precision in function identification and handling during the migration process.
C.2 Challenge 2
Figure 9: A migration case for challenge 2. In this migration case $s_{\text{old}}=s_{\text{new}}$. The green code denotes modifications originating from the out-of-tree kernel patch.
In the migration case shown in Figure 9, we need to locate the target code segment $s_{\text{new}}$, which encompasses lines 4 to 10, within the code file $\text{file}_{\text{new}}$ of the new Linux kernel version. However, when LLMs perform this task, they only retrieve the code segment from line 7 to line 10. As a result, the migrated custom module code exhibits deficiencies due to the missing lines. This issue underscores the limitations of LLMs in accurately identifying precise code segments, suggesting a need for enhanced alignment strategies to improve the reliability of migration tasks.
C.3 Challenge 3
Figure 10: A migration case for challenge 3. (a) The legacy Linux kernel code snippet $s_{\text{old}}$. (b) The updated Linux kernel code snippet $s_{\text{new}}$. The green code denotes modifications originating from the out-of-tree kernel patch.
As shown in Figure 10, in the legacy Linux kernel code snippet $s_{\text{old}}$, the function ptep_get_and_clear is defined. In the updated Linux kernel code snippet $s_{\text{new}}$, this function has been decomposed into two separate definitions: __ptep_get_and_clear and ptep_get_and_clear. The modifications introduced by our out-of-tree kernel patch are located within the definition of the __ptep_get_and_clear function in the $s_{\text{new}}$ code snippet. When employing LLMs directly to retrieve $s_{\text{new}}$ from $\text{file}_{\text{new}}$, the LLMs tend to overlook the definition of __ptep_get_and_clear, focusing instead on the definition of ptep_get_and_clear present in the new version code. Consequently, during the subsequent phase of migrating the out-of-tree kernel patch, the correct migration point cannot be identified, leading to erroneous migration. This issue highlights the difficulties LLMs face in handling the fragmentation of code during version updates, indicating a need for improved methods to accurately locate and integrate all relevant code fragments for successful migration
C.4 Challenge 4
Figure 11: A migration case for challenge 4. The green code denotes modifications originating from the out-of-tree kernel patch.
As shown in Figure 11, to accurately obtain the migrated out-of-tree kernel patch code $s^{\prime}_{\text{new}}$, it is essential to perform two modifications on the new Linux kernel code segment $s_{\text{new}}$ (specifically, adding the code snippet #ifdef CONFIG_HIVE at two locations). However, when LLMs undertake this task, they either misidentify the migration positions or only execute one of the required modifications. This results in the failure of the out-of-tree kernel patch code migration. This issue reveals the limitations of LLMs in interpreting the precise context required for accurate migration, suggesting a need for more refined techniques to enhance the models’ ability to infer migration points based on the given information correctly.
Appendix D MigGPT Modules
D.1 Examples of CFP
Figure 12 illustrates a segment of code alongside its corresponding CFP. The CFP sub-statement in the second row of Figure 12 (b), IfdefNode, represents the second line of the code snippet in Figure 12 (a). This indicates an #ifdef statement that spans from line 2 to line 4 (pos=2, end=4) of the code segment, with the critical identifier being ARM_64_SWAPPER_USES_MAPS (name=’ARM_64_SWAPPER_USES_MAPS’).
Figure 12: (a) A code snippet. (b) Corresponding CFP of the code snippet.
D.2 Examples of Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3.
For challenge 1, we construct a “Namespace Prompt” to specify the signatures of the code snippet $s_{\text{old}}$. By constructing the code fingerprint structure $\text{CFP}_{\text{old}}$ from $s_{\text{old}}$ as shown in Figure 8, we can extract FuncDef statements that contain the code signatures (Figure 13), thereby generating a “Namespace Prompt” that describes these signatures. Consequently, the LLM will focus its attention on the function definition __pte_free_tlb rather than on the similar function definition __pmd_free_tlb. This Namespace prompt enhances the LLM’s ability by providing a precise description of the target code, allowing the LLM to focus more accurately on the relevant code snippet and improving the precision of the retrieval.
Figure 13: The CFP statement on line 10 of Figure 8
For challenge 3, we extract the associated function calls of the code snippet to provide comprehensive code context. As shown in Figure 10 (b), when retrieving $s_{\text{new}}$, the LLM can only find the definition snippet of the function ptep_get_and_clear (lines 16-21) and overlooks the definition snippet of the internally called function __ptep_get_and_clear (lines 1 to 14). To address this challenge, it is necessary to supplement the initially retrieved $s_{\text{tmp}}$ from $\text{file}_{\text{new}}$ with its invoked associated functions, ultimately obtaining a complete code snippet $s_{\text{new}}$. It should be noted that the function ptep_get_and_clear often invokes many functions (such as contpte_try_unfold on line 19), which also appear in $s_{\text{old}}$ (line 4 of Figure 10 (a)) and are not what we require. Therefore, we need to select only those associated functions that are invoked within $s_{\text{tmp}}$ but not by $s_{\text{old}}$ to form the complete code snippet $s_{\text{new}}$.
D.3 Examples of Migration Augmentation Module
The migration augmentation module is primarily designed to address challenge 4. Specifically, as shown in Figure 11, we conduct a comparative analysis between the code fingerprint structures $\text{CFP}_{\text{old}}$ and $\text{CFP}^{\prime}_{\text{old}}$ of the code snippets to ascertain that there are two primary migration points. The first point is located after the comment statement Tial call offset... and before the macro definition statement #define PROLOGUE_OFFSET.... The second point is situated after the statement const struct bpf_prog... and before the statement const int idx0=ctx->idx. By constructing the “Location Prompt”, we enable the LLM to precisely locate the migration points, thereby successfully completing the task of migrating and maintaining the out-of-tree kernel patch.
D.4 Prompts
[…]
Figure 14: The prompts of MigGPT.
Figure 15: Comparison of Consistency with Human Match at Different Thresholds for CodeBLEU.
Appendix E Settings
E.1 Line Edit Distance
The line edit distance is a measure of the difference between two code snippets. It is defined as the minimum number of single-line edit operations (insertions, deletions, or substitutions) required to transform one line into another.
Given two code snippets $A=\{a_{i}\}_{i=1}^{n}$ and $B=\{b_{j}\}_{j=1}^{m}$ with line lengths $|A|=n$ and $|B|=m$, the line edit distance $D(A,B)$ can be defined recursively as follows:
$D(A,B)=\begin{cases}\max(n,m)&\text{if }\min(n,m)=0,\\ \min\begin{cases}D(\text{prefix}(A,n-1),B)+1,&\text{(deletion)}\\ D(A,\text{prefix}(B,m-1))+1,&\text{(insertion)}\\ D(\text{prefix}(A,n-1),\text{prefix}(B,m-1))+\mathbb{I}(a_{n}\neq b_{m})&\text{(substitution)}\end{cases}&\text{otherwise.}\end{cases}$
Where:
  1. 1.
    $\text{prefix}(A,k)=\{a_{i}\}_{i=1}^{k}$ denotes the first $k$ lines of code snippet $A$.
  2. 2.
    $\mathbb{I}(a_{i}\neq b_{j})$ is an indicator function that equals 1 if $a_{i}\neq b_{j}$ and 0 otherwise.
  3. 3.
    The three cases in the recursion correspond to:
    1) Deletion: Delete the last line of $A$ and compute $D(\text{prefix}(A,n-1),B)$.
       2) Insertion: Insert the last line of $B$ into $A$ and compute $D(A,\text{prefix}(B,m-1))$.
       3) Substitution: Replace the last line of $A$ with the last line of $B$ (if they differ) and compute $D(\text{prefix}(A,n-1),\text{prefix}(B,m-1))$.
E.2 Threshold of CodeBLEU
CodeBLEU (Ren et al., 2020) is an automated metric designed to evaluate the quality of code generation, specifically tailored for tasks involving the generation of programming code. By integrating both syntactic and semantic features of code, CodeBLEU provides a similarity score ($[0,1]$) between two code snippets. We employ CodeBLEU as a measure of “semantic match” and investigate the alignment between CodeBLEU-based “semantic matches” and “human matches” across various thresholds. As illustrated in Figure 15 and Table 7, we identify a threshold of 0.9 as optimal for our proposed benchmark, ensuring a high degree of consistency between “semantic matches” derived from CodeBLEU and those determined by human evaluation.
Table 7: The results of MigGPT, compared to the ground truth, are presented in terms of the number of correct examples under both CodeBLEU “semantic match” and “human match”. Here, “CodeBLEU-0.8” denotes a CodeBLEU classification threshold set at 0.8.
Metric Type GPT-4-turbo DeepSeek-V2.5 DeepSeek-V3 Average Retrieval Migration Retrieval Migration Retrieval Migration Retrieval Migration Human Match Type1 77 68 77 64 76 70 77 67 Type2 49 38 50 39 49 41 49 39 All 126 106 127 103 125 111 126 107 CodeBLEU-0.8 Type1 78 77 79 77 77 77 78 77 Type2 46 45 50 48 51 48 49 47 All 124 122 129 125 128 125 127 124 CodeBLEU-0.85 Type1 78 69 79 70 76 74 78 71 Type2 46 43 50 46 51 45 49 45 All 124 112 129 116 127 119 127 116 CodeBLEU-0.9 Type1 76 66 78 64 75 71 76 67 Type2 46 42 50 41 51 43 49 42 All 122 108 128 105 126 114 125 109 CodeBLEU-0.95 Type1 76 62 78 57 75 67 76 62 Type2 45 40 49 38 49 41 48 40 All 121 102 127 95 124 108 124 102
E.3 Variant of MigGPT
We implement four variants for the ablation study:
  1. 1.
    MigGPT-No-Retrieval-Augmentation: the Retrieval Augmentation Module is deactivated, causing no constraint on the namespace of code snippets.
  2. 2.
    MigGPT-No-Retrieval-Alignment: the Retrieval Alignment Module is deactivated, leading to the absence of descriptions for the starting and ending line information of code snippets.
  3. 3.
    MigGPT-No-Migration-Augmentation: the Migration Augmentation Module is disabled. The LLMs will not have the assistance of additional analytical information when completing migration tasks.
  4. 4.
    MigGPT-No-CFP: Replace all components of MigGPT that require CFP participation (including code snippet invocation relationship analysis, anchor function identification, and migration location detection) with implementations utilizing LLMs.